Write the equation of this arithmetic sequence and find the value of the 25th term. -20, -13, -6, 1, 8, 15...
Answer by amarjeeth123(574)

This is an arithmetic progression(AP).
First term a=-20
Common difference d=7
The nth term of an A.P. is a+(n-1)d
Plugging in the values we get,
The 25th term of the given A.P. is -20+(25-1)7=-20+175-7=175-27=148.
Answer=148